# Triad: A Framework Leveraging a Multi-Role LLM-based Agent to Solve Knowledge Base Question Answering

### v1

@@ Abstract
Recent progress with LLM-based agents has shown promising results across various tasks. However, their use in answering questions from knowledge bases remains largely unexplored. Implementing a KBQA system using traditional methods is challenging due to the shortage of task-specific training data and the complexity of creating task-focused model structures. In this paper, we present Triad, a unified framework that utilizes an LLM-based agent with three roles for KBQA tasks. The agent is assigned three roles to tackle different KBQA subtasks: agent as a generalist for mastering various subtasks, as a decision maker for the selection of candidates, and as an advisor for answering questions with knowledge. Our KBQA framework is executed in four phases, involving the collaboration of the agent’s multiple roles. We evaluated the performance of our framework using three benchmark datasets, and the results show that our framework outperforms state-of-the-art systems on the LC-QuAD and YAGO-QA benchmarks, yielding F1 scores of 11.8% and 20.7%, respectively.
@@ 1 Introduction
A question-answering system is designed to extract information by converting a natural language question into a structured query that can retrieve precise information from an existing knowledge base Omar et al. (2023a). The resolution of Knowledge Base Question Answering (KBQA) typically involves phases including question understanding, URI linking, and query execution. Traditional KBQA systems require the use of specialized models trained with domain datasets for question parsing and entity linking Hu et al. (2018); Omar et al. (2023a); Hu et al. (2021). Large language models (LLMs), however, have shown promising competencies in in-context learning using task-specific demonstrations Dong et al. (2022). LLMs have recently been employed as agents in the execution of complex problems. A framework that employs LLM-augmented agents can generate actions or coordinate multiple agents, thus improving the capacity to handle complex situations Liu et al. (2023). Despite the remarkable performance of LLMs in various tasks as evidenced in previous studies, a comprehensive qualitative and quantitative evaluation of KBQA frameworks empowered with an LLM-based agent remains a challenge.
@@
Recent research on KBQA with LLMs has focused primarily on highlighting the inability of LLMs to generate complete factoid results Hu et al. (2023b); Tan et al. (2023c) or demonstrating their potential efficacy in future research Omar et al. (2023b); Tan et al. (2023b). Other studies on LLM-based KBQA concentrates on generating answers by prompt learning and incorporating external knowledge bases Baek et al. (2023); Tan et al. (2023a). Concurrently, LLMs can be deployed to address subtasks within a complex task. In the Text2SQL challenge, the work Li et al. (2023a, b) attempts to utilize an LLM with graph-aware structural inductive bias or as a database interface to handle minor issues. In another recent studyDong et al. (2023), five distinct roles of mathematicians, enacted by an LLM, are introduced to collaboratively prove a theorem. Such research has spurred our exploration into the following question: Whether an LLM-based agent can solve KBQA tasks by serving as multiple roles.
@@
In this study, we introduce Triad, a unified framework that leverages an LLM-based agent with three roles to address KBQA tasks. Specifically, we implement the agent consisting of an LLM as the core, supplemented by various task-specific modules such as memory and executing functions. The agent is assigned three distinct roles: a generalist (G-Agent) adept at mastering numerous small tasks by the given examples, a decision maker (D-Agent) proficient at identifying options and selecting candidates, and an advisor (A-Agent) skilled at providing answers using internal and external knowledge. The cooperation of these agent roles composes a KBQA process containing four phases: question parsing, URI linking, query construction, and answer generation. We evaluate our framework on three benchmark datasets in various difficulties. The results show that our framework outperforms state-of-the-art systems, demonstrated by 11.8% and 20.7% F1 scores on the LC-QuAD and YAGO-QA benchmarks, respectively. The contributions of this study can be summarized as follows.
  • We propose Triad, the first framework that leverages an LLM-based agent to solve KBQA tasks in all its four phases, without specialized training models.
  • We implement an LLM-based agent with various task-specific modules that can act as three roles, including a generalist, a decision maker, and an advisor, to collaboratively solve subtasks in KBQA.
  • We evaluate the performance of our multi-role agent framework. The results show a competitive ability compared to both state-of-the-art KBQA systems and pure LLM methods [^1] Code and datasets used to achieve the performance will be made public if the paper is accepted..
@@ 2.2 Roles of LLM-based Agent
We employ an LLM-based agent to solve a series of subtasks by assuming three distinct roles, which collaboratively execute the four phases of KBQA.
@@
(G-Agent) is capable of mastering various tasks, given a few examples.
@@
(D-Agent) concentrates on the identification of options and the selection of candidates.
@@
(A-Agent) is skilled in answering questions with the aid of both external and internal knowledge.
@@ 3.3 A-Agent as a Comprehensive Advisor
4 Experimental Evaluation
4.1 Indexed Knowledge Bases
@@
4.2 KBQA Benchmark Datasets
We evaluate our framework on datasets including YAGO-QA, LC-QuAD 1.0, and QALD-9, which have various difficulties in interpreting the questions. These datasets contain questions in English, paired with their respective SPARQL queries, and accurate responses derived from a specific knowledge base. QALD-9 Usbeck et al. (2018) and LC-QuAD 1.0 Trivedi et al. (2017) are frequently used to evaluate QA systems in different versions of DBpedia. The recently published YAGO-QA in Omar et al. (2023a), features questions accompanied by annotated SPARQL queries and answers sourced from the YAGO. The statistics for three benchmarks, along with their associated knowledge bases, are depicted in Table 1.
@@ 4.2 KBQA Benchmark Datasets
4.3 Baseline Methods
4.4 Implementation Details
Triad is implemented with Python 3.9. We incorporate LLM capabilities to our multi-role agent via OpenAI’s API services. The names of entities and relations from knowledge bases are indexed in an ElasticSearch 7.5.2 server for text matching. All triples are imported into an SPARQL endpoint of Virtuoso 07.20.3237 for traversal and retrieval. Triad requires four hyperparameters: the number of examples G-Agent uses for subtask learning, the number of candidates D-Agent selects for entity and relation linking, and the retry times for handling non-response SPARQLs. The optimal values for these parameters are 3, 2, 2, and 3, respectively. The framework and its variants are tested five times on each benchmark, with the average scores reported as the final results. For traditional systems, we report the results recorded in their papers. For pure LLM baselines, we write prompts to hire an LLM to answer questions directly referring to examples, and then link the mentions from the responses to the URIs in our indexed knowledge bases via built-in similarity search.
4.5 Performance Comparison
Table 2: The performance of our Triad on three benchmarks, comparing with traditional KBQA systems (full-shot) and pure LLM (few-shot) baselines. The optimal and suboptimal scores are highlighted with bold and underlined text, respectively.
@@ 4.5 Performance Comparison
4.6 Agent Role Ablation Study
@@
Table 3: Ablation study on the roles of LLM-based agent by eliminating an element or downgrading the capability.
@@
4.7 Role Hyperparameter Analysis
We concentrate on three hyperparameters of roles, including the number of examples ($\mathcal{N}$) provided for G-Agent to learn subtasks, the number of URI candidates ($\mathcal{K}$) selected by D-Agent for query construction, and the number of retry times ($\mathcal{T}$) launched by A-Agent when there is no response. The space of hyperparameters is given as follows:
  • Example numbers $\mathcal{N}$: {1,2,3}
  • URI candidate numbers of entities and relations $\mathcal{K}$: {(1,1),(1,2),(2,2),(2,3)}
  • Retry times $\mathcal{T}$: {1,2,3}
Table 4 presents the F1 results of Triad’s performance, employing three hyperparameters on two benchmarks. From the results, we discover that:
@@ 4.7 Role Hyperparameter Analysis
More examples provided to G-Agent do not always improve the performance. The efficacy of G-Agent is significantly influenced by the quality of the provided examples.
@@
Persistently attempting to construct and execute SPARQL queries, despite the initial lack of results, is an effective strategy that improves the probability of obtaining accurate answers. Considering the efficiency of overall execution, we set the maximum retry times as 3 in practice.
4.8 Linking Survival Analysis
The process of linking is a relatively complex subtask in both the Text2SQL and the KBQA process Li et al. (2023b). Calculating the recall ratio of accurate URIs using D-Agent provides clarity on which step most adversely impacts performance. In the entity linking phase, considering all URIs of entities in the testing set as the ground truth of the linking results, 80. 75% of the correct URIs are contained from the output of the entity matching filter in D-Agent and 70. 50% of the correct URIs are retained from the entity selection performed by the LLM in D-Agent. Whereas, in the relation linking phase, only 52. 54% of the correct relation URIs survive from the selection of LLM, which indicates a greater difficulty in relation linking.
4.9 Complex Case Analysis
@@ 5.1 SPARQL-based and LLM-based KBQA
Traditional SPARQL-based KBQA methods transform natural language queries into SPARQL requests for data extraction. Various specific models are employed either for question understanding or URI linking, utilizing domain-based training datasets. Hu et al. (2018) introduces a semantic query graph to structurally represent the natural language query, thereby simplifying a SPARQL-based KBQA into a subgraph matching problem. Hu et al. (2021) proposes an entity description graph to represent natural language queries for question parsing and element linking. The most recent system Omar et al. (2023a) restructures the question parsing task as a text generation issue using a sequence-to-sequence model, as opposed to a rule-based transformation. With the advent of large language models (LLMs), certain phases of KBQA can be enhanced with LLM-integrated methods. Baek et al. (2023) aims to augment LLM-based QA tasks with pertinent facts extracted from knowledge bases, offering a fully zero-shot architecture. Tan et al. (2023a) leverages the general applicability of LLMs to filter linking candidates by making selections via few-shot in-context learning. Omar et al. (2023b) provides a thorough comparison between LLMs and QA systems in KBQA tasks, recommending further studies to improve the KBQA process with LLM capabilities. However, apart from the above studies, our investigation proposes a unified framework for KBQA tasks, which incorporates both an LLM and few-shot or zero-shot prompting from a systematic perspective.
@@ 5.2 LLM-based Agents for Complex Tasks
LLMs have recently gained significant attention due to their ability to approximate human-level intelligence through the acquisition of vast amounts of web knowledge. This has led to numerous studies focusing on LLM-based agents. A recent surveyWang et al. (2023) proposes a unified architectural design for LLM-based agents, which consists of four modules that include profile, memory, plan, and action. CHATDBHu et al. (2023a) employs an LLM controller to generate SQL instructions, enabling it to manipulate databases, which allows for symbolic memory and complex multi-hop reasoning. ARTParanjape et al. (2023) uses a frozen LLM to generate intermediate reasoning steps and further integrates tools for new tasks with minimal human intervention. ToolformerSchick et al. (2023) takes a different approach by training an LLM to plan and execute tools for the next token prediction by learning API calls generation. ReActYao et al. (2023) focuses on overcoming LLM hallucination by interacting with external knowledge bases, thus generating interpretable task-solving strategies. Divergent from the aforementioned research, our framework concentrates on the resolution of KBQA tasks by introducing a multi-role LLM-based agent that specializes in various types of subtask.
@@ 6 Conclusion
In this study, we aim to bridge the gap between KBQA tasks and the investigation of LLM-based agents. We introduce Triad, a unified framework to address the KBQA task through an LLM-based agent acting as three roles. This includes a generalist capable of mastering diverse tasks given minimal examples, a decision-maker concentrating on option identification and candidate selection, and an advisor skilled in answering questions with the aid of both external and internal knowledge. Triad achieves the best or competitive performance across three benchmark datasets compared to traditional KBQA systems trained on full-size QA data and pure LLM models. In future research, we plan to broaden our framework to handle more intricate scenarios, such as accommodating advanced queries and multi-hop implicit reasoning.
@@
All datasets utilized in this study are publicly available and we have adhered to ethical considerations by not introducing additional information as input during LLM training and LLM text generation.

### v3

@@ Abstract
Recent progress with LLM-based agents has shown promising results across various tasks. However, their use in answering questions from knowledge bases remains largely unexplored. Implementing a KBQA system using traditional methods is challenging due to the shortage of task-specific training data and the complexity of creating task-focused model structures. In this paper, we present Triad, a unified framework that utilizes an LLM-based agent with multiple roles for KBQA tasks. The agent is assigned three roles to tackle different KBQA subtasks: agent as a generalist for mastering various subtasks, as a decision maker for the selection of candidates, and as an advisor for answering questions with knowledge. Our KBQA framework is executed in four phases, involving the collaboration of the agent’s multiple roles. We evaluated the performance of our framework using three benchmark datasets, and the results show that our framework outperforms state-of-the-art systems on the LC-QuAD and YAGO-QA benchmarks, yielding F1 scores of 11.8% and 20.7%, respectively.
@@ 1 Introduction
A question-answering system is designed to extract information by converting a natural language question into a structured query that can retrieve precise information from an existing knowledge base Omar et al. (2023a). The resolution of Knowledge Base Question Answering (KBQA) typically involves phases including question understanding, URI linking, and query execution. Traditional KBQA systems require the use of specialized models trained with domain datasets for question parsing and entity linking Hu et al. (2018); Omar et al. (2023a); Hu et al. (2021). Large language models (LLMs), however, have shown promising competencies in in-context learning using task-specific demonstrations Dong et al. (2022). LLMs have recently been employed as agents in the execution of complex problems. A framework that employs LLM-augmented agents can generate actions or coordinate multiple agents, thus improving the capacity to handle complex situations Liu et al. (2023). Despite the remarkable performance of LLMs in various tasks as evidenced in previous studies, a comprehensive qualitative and quantitative evaluation of KBQA frameworks empowered with an LLM-based agent remains insufficiently explored.
Figure 1: A system with multiple roles who focus on sub-problems of each phase to solve a complex task.
@@
Studies on KBQA with LLMs has attracted considerable attention. Some works focus primarily on highlighting the inability of LLMs to generate complete factoid results Hu et al. (2023b); Tan et al. (2023c) or demonstrating their potential efficacy in future research Omar et al. (2023b); Tan et al. (2023b). Other works concentrates on generating answers by prompt learning and incorporating external knowledge bases Baek et al. (2023); Tan et al. (2023a). Concurrently, LLMs can be deployed to address each phase within Text2SQL challengesLi et al. (2023, 2024) or theorem proof tasksDong et al. (2023). However, each phase of KBQA can be further decomposed into subtasks and completed through an agentic approach that provides feedback and cooperation. Additionally, decomposing the task reduces the complexity of cooperative working by allowing each role to concentrate on smaller sub-problemsWang et al. (2020). As illustrated in Figure 1, three roles in an organization work together to provide the final answer for the overall task. The above observations spur our exploration into the following question: Whether an LLM-based agent can solve KBQA tasks by serving as multiple roles.
@@
In this study, we introduce Triad, a unified framework that leverages an LLM-based agent with three roles to address KBQA tasks. Specifically, we implement the agent consisting of an LLM as the core, supplemented by various task-specific modules such as memory and executing functions. The agent is assigned three distinct roles: a generalist (G-Agent) adept at mastering numerous small tasks by the given examples, a decision maker (D-Agent) proficient at identifying options and selecting candidates, and an advisor (A-Agent) skilled at providing answers using internal and external knowledge. The cooperation of these agent roles composes a KBQA process containing four phases: question parsing, URI linking, query construction, and answer generation. We evaluate our framework on three benchmark datasets in various difficulties. The results show that our framework outperforms state-of-the-art systems, demonstrated by 11.8% and 20.7% F1 scores on the LC-QuAD and YAGO-QA benchmarks, respectively.
The contributions of this study can be summarized as follows: (1) We propose Triad, the first framework that leverages an LLM-based agent to solve KBQA tasks in all its four phases, without specialized training models. (2) We implement an LLM-based agent with various task-specific modules that can act as three roles, including a generalist, a decision maker, and an advisor, to collaboratively solve KBQA via focusing on subtasks. (3) We evaluate the performance of Triad. The results show a competitive ability compared to both state-of-the-art KBQA systems and pure LLM methods [^1] Code and datasets used to achieve the performance will be made public if the paper is accepted..
@@ 2.2 Roles of LLM-based Agent
Drawing an analogy to a software development scenario, where coders complete small development tasks, with the process and plan being decided by the manager, and ultimately the outcome inspected by the leader, we assign the following three roles to an LLM-based agent to solve the KBQA task:
@@
(G-Agent) is capable of mastering various small tasks by providing a few examples.
@@
(D-Agent) adepts at analyzing options and providing candidate results as procedural feedback.
@@
(A-Agent) is skilled in providing final answers with the aid of both external and its own knowledge.
@@ 3.3 A-Agent as a Comprehensive Advisor
4 Performance Evaluation
4.1 Experimental Settings
Indexed Knowledge Bases:
@@ 4.1 Experimental Settings
KBQA Benchmark Datasets:
We evaluate our framework on datasets including YAGO-QA, LC-QuAD 1.0, and QALD-9, which have various difficulties in interpreting the questions. These datasets contain questions in English, paired with their respective SPARQL queries, and accurate responses derived from a specific knowledge base. QALD-9 Usbeck et al. (2018) and LC-QuAD 1.0 Trivedi et al. (2017) are frequently used to evaluate QA systems with DBpedia. The recently published YAGO-QA in Omar et al. (2023a), features questions accompanied by annotated SPARQL queries sourced from YAGO. The statistics for three benchmarks, along with their associated knowledge bases, are depicted in Table 1.
@@
Table 2: The performance of our proposed Triad on three benchmarks, comparing with traditional KBQA systems (full-shot) and pure LLM (few-shot) baselines. The optimal and suboptimal scores are highlighted with bold and underlined text, respectively.
@@
Baseline Methods:
Implementation Details:
Triad is implemented with Python 3.9. We incorporate LLM capabilities to our multi-role agent via OpenAI’s API services. The names of entities and relations from knowledge bases are indexed in an ElasticSearch 7.5.2 server for text matching. All triples are imported into an SPARQL endpoint of Virtuoso 07.20.3237 for retrieval. Triad requires four hyperparameters: the number of examples G-Agent uses for subtask learning, the number of candidates D-Agent selects for entity and relation linking, and the retry times for handling non-response SPARQLs. The optimal values for these parameters are 3, 2, 2, and 3, respectively. The framework and its variants are tested five times on each benchmark, with the average scores reported as the final results. For traditional systems, we report the results recorded in their papers. For pure LLM baselines, we write prompts to hire an LLM to answer questions directly referring to examples, and then link the mentions from the responses to the URIs in our indexed knowledge bases via built-in similarity search.
4.2 Performance Comparison
@@
4.3 Study on Capabilities of Agent Roles
@@
Table 3: Study on the roles of LLM-based agent by eliminating an element or downgrading the capability.
@@
4.4 Analysis of Role Hyperparameters
We concentrate on three hyperparameters of roles, including the number of examples ($\mathcal{N}\in\{1,2,3\}$) provided for G-Agent to learn subtasks, the number of URI candidates ($\mathcal{K}\in\{(1,1),(1,2),(2,2),(2,3)\}$) selected by D-Agent for query construction, and the number of retry times ($\mathcal{T}\in\{1,2,3\}$) launched by A-Agent when there is no response. Table 4 presents the F1 results of Triad’s performance, employing three hyperparameters on two benchmarks. We discover that:
@@ 4.4 Analysis of Role Hyperparameters
More examples provided to G-Agent do not always improve the performance. The efficacy of G-Agent is significantly influenced by the quality of examples.
@@
Persistently attempting to construct and execute SPARQL queries is an effective strategy that improves the probability of obtaining accurate answers. Considering the efficiency of overall execution, we set the maximum retry times as 3 in practice.
4.5 Analysis of Linking Recall
The process of linking is a relatively complex subtask in both the Text2SQL and the KBQA process Li et al. (2024). Calculating the recall ratio of accurate URIs using D-Agent provides clarity on which step most adversely impacts performance. In the entity linking phase, considering all URIs of entities in the testing set as the ground truth of the linking results, 80. 75% of the correct URIs are contained from the output of the entity matching filter in D-Agent and 70. 50% of the correct URIs are retained from the entity selection performed by the LLM in D-Agent. Whereas, in the relation linking phase, only 52. 54% of the correct relation URIs survive from the selection of LLM, which indicates a greater difficulty in relation linking.
4.6 Study on Complex Cases
@@ 5.1 SPARQL-based and LLM-based KBQA
Traditional KBQA methods transform natural language queries into SPARQL requests for data extraction. Specific models are employed either for question understanding or URI linking, utilizing domain-based training datasets. Hu et al. (2018) introduces a semantic query graph to structurally represent the natural language query, thereby simplifying the task into a subgraph matching problem. Hu et al. (2021) proposes an entity description graph to represent natural language queries for question parsing and element linking. Omar et al. (2023a) restructures the question parsing task as a text generation issue using a sequence-to-sequence model. With the advent of LLMs, certain phases of KBQA can be enhanced with LLM-integrated methods. Baek et al. (2023) aims to augment LLM-based QA tasks with pertinent facts extracted from knowledge bases, offering a fully zero-shot architecture. Tan et al. (2023a) leverages the general applicability of LLMs to filter linking candidates by making selections via few-shot in-context learning. Omar et al. (2023b) provides a thorough comparison between LLMs and QA systems, recommending further studies to improve KBQA with LLM capabilities. However, apart from the above studies, our study proposes a complete framework incorporating both an LLM and few-shot learning across all KBQA phases from a systematic perspective.
@@ 5.2 LLM-based Agents for Complex Tasks
LLMs have recently gained significant attention due to their ability to approximate human-level intelligence through the acquisition of vast amounts of web knowledge. This has led to numerous studies focusing on LLM-based agents. A recent surveyWang et al. (2023) proposes a unified architectural design for LLM-based agents, which consists of four modules that include profile, memory, plan, and action. CHATDBHu et al. (2023a) employs an LLM controller to generate SQL instructions, enabling it to manipulate databases, which allows for symbolic memory and complex multi-hop reasoning. ARTParanjape et al. (2023) uses a frozen LLM to generate intermediate reasoning steps and further integrates tools for new tasks with minimal human intervention. ToolformerSchick et al. (2024) takes a different approach by training an LLM to plan and execute tools for the next token prediction by learning API calls generation. ReActYao et al. (2023) focuses on overcoming LLM hallucination by interacting with external knowledge bases, thus generating interpretable task-solving strategies. Divergent from the aforementioned studies, our framework concentrates on the solving KBQA tasks by introducing a multi-role LLM-based agent that specializes in various subtasks distributed across different phases.
@@ 6 Conclusion
In this study, we aim to bridge the gap between KBQA tasks and the investigation of LLM-based agents. We introduce Triad, a framework to address the KBQA task through an LLM-based agent acting as multiple roles, including a generalist capable of mastering diverse tasks given minimal examples, a decision-maker concentrating on option analysis and candidate selection, and an advisor skilled in answering questions with the aid of both external and internal knowledge. Triad achieves the best or competitive performance across three benchmark datasets compared to traditional KBQA systems and pure LLM models. In future research, we plan to broaden our framework to handle more intricate questions, such as multi-hop reasoning, and exploring the integration between our framework and retrieval-augmented generation.
@@
All datasets utilized in this study are publicly available and we have adhered to ethical considerations by not introducing additional information as input during LLM text generation.